ABC Enterprises has five jobs that must be scheduled for production at a work center. Below are the processing times and due dates for each job.

                                   

Job

Processing Time
(in days)

Due Date
(days from today)

1

7

14

2

5

21

3

3

10

4

2

7

5

6

9

                                                                                                                          

Schedule the jobs by sequencing them using the Earliest Due Date (EDD) priority rule.
What is the average flow time? Show your Work and round your answer to one decimal point.

Job

Processing Time

(in days)

Due Date
(days from today)

Flow Time

                                               

Schedule the jobs by sequencing them using the Shortest Processing Time (SPT) priority rule.
What is the average flow time? Show your Work and round your answer to one decimal point.

                                               

Job

Processing Time

(in days)

Due Date
(days from today)

Flow Time
Answer #1

EDD

JOB PROCESSING TIME DUE TIME FLOW TIME JOB LATENESS
D 2 7 0 + 2 = 2
E 6 9 2 + 6 = 8
C 3 10 8 + 3 = 11 1
A 7 14 11 + 7 = 18 4
B 5 21 18 + 5 = 23 2

TOTAL FLOW TIME = 2 + 8 + 11 + 18 + 23 = 62
AVERAGE COMPLETION TIME = FLOW TIME / N = 62 / 5 = 12.4

SPT

JOB PROCESSING TIME DUE TIME FLOW TIME JOB LATENESS
D 2 7 0 + 2 = 2
C 3 10 2 + 3 = 5
B 5 21 5 + 5 = 10
E 6 9 10 + 6 = 16 7
A 7 14 16 + 7 = 23 9

TOTAL FLOW TIME = 2 + 5 + 10 + 16 + 23 = 56
AVERAGE COMPLETION TIME = FLOW TIME / N = 56 / 5 = 11.2

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.